# Adaptive Distribution Masked Autoencoders for Continual Test-Time Adaptation

### v1

@@ A.2 Additional Ablation Study
Components Effectiveness on ImageNet-to-ImageNet-C. We conduct an additional experiment to evaluate each component of our proposed method on the ImageNet-to-ImageNet-C CTTA task. Consistent with our submission, we perform four sets of ablation studies. As shown in Table 8, the first set of experiments (Ex1) is applying random masking strategy to establish consistency constraints between the model outputs generated from the masked target samples and those from the original target samples. This obtains a 6.4% reduction in the error rate in contrast to the source method (Ex0). Secondly, with the implementation of the Distribution-aware Masking (DaM) mechanism, the error rate (Ex2) is further reduced to 47.9%, validating that DaM significantly enhances the model’s ability to understand the target domain distribution. The remaining two sets of experiments (Ex3 and Ex4) are reconstructing the Histogram of Oriented Gradients (HOG) feature representations based on two masking strategies. Random masking strategy with HOG reconstruction scheme (Ex3) achieves a 1.7% reduction in the error rate compared to use random masking strategy individually (Ex1). Our method (Ex4) outperforms others, showcasing the best results, with a remarkable 13.3% reduction in error rate compared to the source method. These results confirm that incorporating HOG reconstruction into the continual adaptation process aids the model in acquiring task-relevant knowledge, particularly in the presence of domain shifts.

### v2

@@ 6 Conclusion
Acknowledgements.
Shanghang Zhang is supported by the National Science and Technology Major Project of China (No. 2022ZD0117801).
@@ A.2 Additional Ablation Study
Components Effectiveness on ImageNet-to-ImageNet-C. We conduct an additional experiment to evaluate each component of our proposed method on the ImageNet-to-ImageNet-C CTTA task. Consistent with our submission, we perform four sets of ablation studies. As shown in Table 8, the first set of experiments (Ex1) is applying random masking strategy to establish consistency constraints between the model outputs generated from the masked target samples and those from the original target samples. This obtains a 6.4% reduction in the error rate in contrast to the source method (Ex0). Secondly, with the implementation of the Distribution-aware Masking (DaM) mechanism, the error rate (Ex2) is further reduced to 47.9%, validating that DaM significantly enhances the model’s ability to understand the target domain distribution. The remaining two sets of experiments (Ex3 and Ex4) are reconstructing the Histogram of Oriented Gradients (HOG) feature representations based on two masking strategies. Random masking strategy with HOG reconstruction scheme (Ex3) achieves a 1.7% reduction in the error rate compared to use random masking strategy individually (Ex1). Our method (Ex4) outperforms others, showcasing the best results, with a remarkable 12.2% reduction in error rate compared to the source method. These results confirm that incorporating HOG reconstruction into the continual adaptation process aids the model in acquiring task-relevant knowledge, particularly in the presence of domain shifts.